OpenStudy (anonymous):

Is the function linear or nonlinear? x −3 −1 0 1 3 y −5 −1 1 3 7 A. linear B. nonlinear

OpenStudy (elonasushchik):

I think nonlinear. Im sure

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

correct

OpenStudy (michele_laino):

please a linear function between the variable x and y is given by the subsequent formula: \[y = ax + b\] where a and b are two parameters

OpenStudy (michele_laino):

for example substituting the pair (0,1) I get: 1 = a*0 + b so b=1 whereas substituting the pair (1,3), I get: 3 = a+b 3 = a + 1 and a=2 so we got a= 2 and b=1 and: \[y = 2x + 1\]

OpenStudy (michele_laino):

so your function is a linear function, please check that all of your ordered pairs make that formula as an identity